$SPAD/src/input intef.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool intef.output
)set message test on
)set message auto off
)clear all
--S 1 of 16
(a*x+b) / (b**2 * x * log(x)**2 + 2*a*b*x**2*log(x) + a**2*x**3 + x)
--R
--R
--R                        a x + b
--R   (1)  --------------------------------------
--R         2        2         2          2 3
--R        b x log(x)  + 2a b x log(x) + a x  + x
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 16
integrate(%,x)
--R
--R
--R   (2)  atan(b log(x) + a x)
--R                                         Type: Union(Expression(Integer),...)
--E 2
--S 3 of 16
((exp(x)-x**2+2*x)/(x**2*(exp(x)+x)**2))*exp((x**2-1)/x+1/(exp(x)+x))
--R
--R
--R                           2       x    3
--R                         (x  - 1)%e  + x
--R                         ----------------
--R                                x    2
--R           x    2           x %e  + x
--R        (%e  - x  + 2x)%e
--R   (3)  ---------------------------------
--R               2   x 2     3  x    4
--R              x (%e )  + 2x %e  + x
--R                                                    Type: Expression(Integer)
--E 3

--S 4 of 16
integrate(%,x)
--R
--R
--R            2       x    3
--R          (x  - 1)%e  + x
--R          ----------------
--R                 x    2
--R             x %e  + x
--R        %e
--R   (4)  ------------------
--R                  x
--R                %e
--R                                         Type: Union(Expression(Integer),...)
--E 4

———-

Here are three which do not have elementary integrals — * —

--S 5 of 16
sin(x)/x
--R
--R
--R        sin(x)
--R   (5)  ------
--R           x
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 16
integrate(%,x)
--R
--R
--R   (6)  Si(x)
--R                                         Type: Union(Expression(Integer),...)
--E 6
--S 7 of 16
x * cot x
--R
--R
--R   (7)  x cot(x)
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 16
integrate(%,x)

--R
--R
--R           x
--R         ++
--R   (8)   |   %J cot(%J)d%J
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E 8
--S 9 of 16
(2 * log(x)**2 - log x - x**2) / (log(x)**3 - x**2 * log x)
--R
--R
--R               2             2
--R        2log(x)  - log(x) - x
--R   (9)  ----------------------
--R                3    2
--R          log(x)  - x log(x)
--R                                                    Type: Expression(Integer)
--E 9
--S 10 of 16
integrate(%,x)
--R
--R
--R         log(log(x) + x) - log(log(x) - x) + 2li(x)
--R   (10)  ------------------------------------------
--R                              2
--R                                         Type: Union(Expression(Integer),...)
--E 10

———-

Real trigonometric functions are handled directly — * —

--S 11 of 16
cos(a * x) / (1 + cos(a * x))
--R
--R
--R           cos(a x)
--R   (11)  ------------
--R         cos(a x) + 1
--R                                                    Type: Expression(Integer)
--E 11
--S 12 of 16
integrate(%,x)
--R
--R

--R         - sin(a x) + a x cos(a x) + a x
--R   (12)  -------------------------------
--R                  a cos(a x) + a
--R                                         Type: Union(Expression(Integer),...)
--E 12
--S 13 of 16
cos(3*x)*sin(2*x)
--R
--R
--R   (13)  cos(3x)sin(2x)
--R                                                    Type: Expression(Integer)
--E 13
--S 14 of 16
integrate(%,x)
--R
--R
--R                  5           3
--R         - 8cos(x)  + 10cos(x)
--R   (14)  ----------------------
--R                    5
--R                                         Type: Union(Expression(Integer),...)
--E 14

———-

Integrals of hyperb. trigs are transformed back into hyperb. trigs.

— * —

--S 15 of 16
cosh(a*x)*sinh(a*x)
--R
--R
--R   (15)  cosh(a x)sinh(a x)
--R                                                    Type: Expression(Integer)
--E 15
--S 16 of 16
integrate(%,x)
--R
--R
--R                  2            2
--R         sinh(a x)  + cosh(a x)
--R   (16)  -----------------------
--R                    4a
--R                                         Type: Union(Expression(Integer),...)
--E 16
)spool
)lisp (bye)

———- 7

References^{[1] nothing}

8